### v1

@@ Abstract
We study the common continual learning setup where an overparameterized model is sequentially fitted to a set of jointly realizable tasks. We analyze the forgetting—loss on previously seen tasks—after $k$ iterations. For linear models, we prove that fitting a task is equivalent to a single stochastic gradient descent (SGD) step on a modified objective. We develop novel last-iterate SGD upper bounds in the realizable least squares setup, and apply them to derive new results for continual learning. Focusing on random orderings over $T$ tasks, we establish universal forgetting rates, whereas existing rates depend on the problem dimensionality or complexity. Specifically, in continual regression with replacement, we improve the best existing rate from $\mathcal{O}((d-\bar{r})/k)$ to $\mathcal{O}(\min(1/\sqrt[4]{k},\sqrt{d-\bar{r}}/k,\sqrt{T\bar{r}}/k))$, where $d$ is the dimensionality and $\bar{r}$ the average task rank. Furthermore, we establish the first rates for random task orderings without replacement. The obtained rate of $\mathcal{O}(\min(1/\sqrt[4]{T},\,(d-r)/T))$ proves for the first time that randomization alone—with no task repetition—can prevent catastrophic forgetting in sufficiently long task sequences. Finally, we prove a similar $\mathcal{O}(1/\sqrt[4]{k})$ universal rate for the forgetting in continual linear classification on separable data. Our universal rates apply for broader projection methods, such as block Kaczmarz and POCS, illuminating their loss convergence under i.i.d. and one-pass orderings.
@@ 1 Introduction
In continual learning, tasks are presented sequentially, one at a time. The goal is for the learner to adapt to the current task—e.g., by fine-tuning using gradient-based algorithms—while retaining knowledge from previous tasks. A central challenge in this setting is termed catastrophic forgetting, where expertise from earlier tasks is lost when adapting to newer ones. Forgetting is influenced by factors such as task similarity and overparameterization (Goldfarb et al., 2024), and is also related to trade-offs like the plasticity-stability dilemma (Mermillod et al., 2013). Continual learning is becoming increasingly important with the rise of foundation models, where retraining is prohibitively expensive and data from prior tasks is often unavailable, e.g., due to privacy concerns or data retention issues.
@@
Previous work has shown, both analytically (Evron et al., 2022, 2023; Kong et al., 2023; Jung et al., 2025; Cai and Diakonikolas, 2025) and empirically (Lesort et al., 2023; Hemati et al., 2024), that forgetting diminishes over time when task ordering is cyclic or random. Different orderings can be explored from multiple perspectives: as a strategy to mitigate forgetting (e.g., by actively ordering an agent’s learning environments); as a naturally occurring phenomenon, such as periodic trends in e-commerce systems; or as a means to model and analyze widely used continual learning benchmarks, such as randomly split datasets.
@@
Our work focuses on a widely studied analytical setting—realizable continual linear regression, where $T$ tasks are learned sequentially over $k$ iterations according to a uniform random ordering. Evron et al. (2022) established that the worst-case expected forgetting lies between $\Omega\left({{1}/{k}}\right)$ and $\mathcal{O}\left({{(d-\bar{r})}/{k}}\right)$, where $d$ is the problem dimensionality, and $\bar{r}$ is the average rank of individual data matrices. This raises a fundamental question with critical implications in highly overparameterized regimes: Does the worst-case forgetting necessarily scale with the dimensionality, and if so, is the dependence indeed linear?
@@
To this end, we bridge continual learning and the literature on last-iterate stochastic gradient descent (SGD) analysis. We revisit an established connection between continual linear regression and the Kaczmarz method for solving systems of linear equations (Kaczmarz, 1937; Evron et al., 2022). Given rank-$1$ tasks, this method is known to perform a normalized stochastic gradient step on the least squares objective, fully minimizing the current task’s loss and implying a “stepwise-optimal” step size. Deepening this connection, we prove that even for general data ranks, learning a task in continual linear regression and performing an update in the Kaczmarz method, are both equivalent to a single SGD step on a modified objective with a constant, stepwise-optimal step size.
@@
Motivated by this, we prove convergence rates for the last iterate of fixed-step size SGD that, crucially, hold for a broad range of step sizes not covered by prior work (e.g., Shamir and Zhang, 2013; Ge et al., 2019; Berthier et al., 2020; Zou et al., 2021; Varre et al., 2021; Wu et al., 2022). Somewhat curiously, prior results either hold only for the average iterate (e.g., Bach and Moulines, 2013), or for small step sizes, bounded away from the stepwise-optimal step size crucial for the continual learning setup (e.g., Varre et al., 2021). We overcome this challenge through a careful combination of analysis techniques for SGD (Srebro et al., 2010; Shamir and Zhang, 2013), further tightening the analysis to accommodate a wider range of step sizes, including the stepwise-optimal one.
@@
Applying our last-iterate analysis to continual regression, we tighten the existing forgetting rate and also establish the first dimension-independent rate (see Table 1). Further, we provide the first rate for random task orderings without replacement, proving that task repetition is not required to guarantee convergence when ${k=T\to\infty}$, thus highlighting the effect of randomization as compared to repetition of tasks. Our results also yield novel rates for the closely related Kaczmarz and Normalized Least Mean Squares methods.
@@
Finally, we prove a matching rate for the squared loss of the broader Projection Onto Convex Sets framework (POCS; see Gubin et al., 1967). This extends our results to continual linear classification on separable data—previously linked to projection algorithms (Evron et al., 2023)—and provides this setting’s first universal rate, independent of the problem’s “complexity”.
@@
To summarize, our main contributions in this paper are as follows:
  • We establish a new reduction from continual linear regression to SGD with a particular choice of a “stepwise-optimal” step size, generalizing ideas from recent work that only applied to rank-$1$ tasks, to tasks of arbitrary rank. This facilitates last iterate-analysis to study forgetting.
@@
  • Our main contribution, building on these techniques, is a set of improved rates of forgetting in continual linear regression, including the first universal rates, independent of the problem dimensionality or complexity, as well as the first rates for without-replacement orderings, indicating task repetition is not mandatory to diminish forgetting. See Table 1 for more details.
@@ 2 Main Setting: Continual Linear Regression
We mainly investigate the fundamental continual linear regression setting, as studied in many theoretical papers. This setting is easy-to-analyze, yet often sheds light on important phenomena in continual learning (e.g., Doan et al., 2021; Evron et al., 2022; Lin et al., 2023; Peng et al., 2023; Goldfarb and Hand, 2023; Li et al., 2023; Goldfarb et al., 2024; Hiratani, 2024).
@@
Formally, the learner has access to a task collection of $T$ finite-dimensional linear regression tasks, i.e., $(\mathbf{X}_{1},\mathbf{y}_{1}),\dots,(\mathbf{X}_{T},\mathbf{y}_{T})$ where $\mathbf{X}_{m}\in\mathbb{R}^{n_{m}\times d},\mathbf{y}_{m}\in\mathbb{R}^{n_{m}}$. For $k$ iterations, the learner learns the tasks according to a task ordering $\tau:\left[{k}\right]\to\left[{T}\right]$. We analyze random orderings, previously studied in continual linear models (e.g., Evron et al., 2022, 2023; Jung et al., 2025).
Definition 1 (Random Task Ordering).
@@
We are now ready to define the learning scheme we study, which, at each iteration, naively minimizes the sum of squared errors for the current regression task.111 This objective is natural for regression tasks; our analysis also extends to the mean squared error (by adjusting $R$ in the results).
@@
   For each task $t=1,\dots,k$:
       $\mathbf{w}_{t}\leftarrow$ Start from $\mathbf{w}_{t-1}$ and minimize the current loss $\mathcal{L}_{\tau(t)}(\mathbf{w})\triangleq{\frac{1}{2}}\left\|{\mathbf{X}_{\tau({t})}\mathbf{w}-\mathbf{y}_{\tau({t})}}\right\|^{2}$           with (S)GD to convergence [^2] Learning to convergence facilitates the analysis, but other analytical choices exist (see Jung et al., 2025).
@@
This scheme was previously linked to the Kaczmarz method and, in a special case, to normalized SGD (Evron et al., 2022). In Section 3, we explain and develop these connections to enable novel analysis.
Our primary assumption is that there exist offline solutions that perfectly solve all $T$ tasks jointly, as assumed in much of the theoretical continual learning literature (e.g., Evron et al., 2022, 2023; Goldfarb et al., 2024; Jung et al., 2025). This assumption simplifies the analysis and rules out cases where forgetting previous tasks is beneficial, as new tasks may directly contradict them. Finally, this assumption is reasonable in highly overparameterized models and is thus linked to the linear dynamics of deep neural networks in the neural tangent kernel (NTK) regime (e.g., see Jacot et al., 2018).
Assumption 1 (Joint Linear Realizability).
@@
Definition 2 (Minimum-Norm Offline Solution).
We specifically denote,
$\mathbf{w}_{\star}\triangleq{\operatorname*{argmin}}_{\mathbf{w}\in\mathcal{W}_{\star}}\left\|{\mathbf{w}}\right\|\,.$
Commonly in continual learning setups, the model performance on past tasks degrades, sometimes significantly, even in linear models (Evron et al., 2022). Our goal is to bound the worst-case degradation, i.e., “forgetting”, in linear models under random task orderings. We follow common definitions (e.g., Doan et al., 2021; Evron et al., 2023), and define forgetting as the average increase in the loss of the last iterate on previous tasks. In realizable settings, forgetting takes the form below.
Definition 3 (Forgetting).
@@
Under arbitrary orderings, Evron et al. (2022) showed forgetting can be “catastrophic” in the sense that $\lim_{k\to\infty}\mathbb{E}\left[F_{\tau}(k)\right]>0$. However, as we show, this cannot happen under the random ordering.
Remark 4 (Forgetting vs. Regret).
While regret and forgetting are related, they can differ significantly (Evron et al., 2022). Regret is a key quantity in online learning, defined as $\frac{1}{2k}\sum_{t=1}^{k}\|{\mathbf{X}_{\tau(t)}\mathbf{w}_{t-1}-\mathbf{y}_{\tau(t)}}\|^{2}$ in our setting. That is, it measures the suboptimality of each iterate on the consecutive task. In contrast, forgetting evaluates an iterate’s performance across all earlier tasks.
@@
Definition 5 (Training Loss).
@@
Our results bound both the expected forgetting and expected training loss, leveraging a key property: expected (in-sample) forgetting can be upper bounded using expected training loss across all tasks. Specifically, Lemma 16 (in Appendix B) states that $\mathbb{E}_{\tau}[F_{\tau}(k)]\leq 2\mathbb{E}_{\tau}\left[\mathcal{L}\left(\mathbf{w}_{k-1}\right)\right]+\frac{\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}}{k}$ in orderings with replacement, where $R\triangleq\max_{m\in\left[{T}\right]}\left\|{\mathbf{X}_{m}}\right\|$ is the data “radius” and the dependence of $\mathbf{w}_{k-1}$ on $\tau_{1},\dots,\tau_{k-1}$ is implicit. Without-replacement orderings yield a related but more refined bound. The additive $\frac{\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}}{k}$ term is negligible compared to the bounds obtained in our work.
@@ 3 Reductions: From Continual Linear Regression to Kaczmarz to SGD
Previous work established connections between continual linear regression and the Kaczmarz method in the realizable case (Evron et al., 2022). We revisit these connections pedagogically to ensure our paper is self-contained. Importantly, this leads to novel links between continual learning, the Kaczmarz method, and SGD on special functions (Schemes 1,2,3), allowing us to improve the rates for continual and Kaczmarz methods by analyzing the last iterate of SGD algorithms instead.
@@
  Input: ${f_{m}(\mathbf{w})=\frac{1}{2}\left\|{\mathbf{X}_{m}^{+}\mathbf{X}_{m}\left({\mathbf{w}-\mathbf{w}_{\star}}\right)}\right\|^{2},\forall m\in\left[{T}\right]}$
@@ 3.1 Revisit: Continual Linear Regression and the Kaczmarz Method
The (block) Kaczmarz method in Scheme 2 (Kaczmarz, 1937; Elfving, 1980) is a classical iterative method for solving a linear system $\mathbf{X}\mathbf{w}=\mathbf{y}$, easily mapped to our learning problem by stacking tasks in blocks, i.e.,
$\mathbf{X}=\begin{pmatrix}\mathbf{X}_{1}\\ \vdots\\ \mathbf{X}_{T}\end{pmatrix}\in\mathbb{R}^{N\times d},\quad\mathbf{y}=\begin{pmatrix}\mathbf{y}_{1}\\ \vdots\\ \mathbf{y}_{T}\end{pmatrix}\in\mathbb{R}^{N},\quad\text{where $N=\sum_{m=1}^{T}n_{m}$.}$
In each iteration, the Kaczmarz method (Scheme 2) perfectly solves the current block, i.e., ${\mathbf{X}_{\tau(t)}\mathbf{w}_{\tau(t)}=\mathbf{y}_{\tau(t)}}$ (to see that, recall that $\mathbf{X}_{\tau(t)}^{+}$ denotes the Moore-Penrose pseudo-inverse of $\mathbf{X}_{\tau(t)}^{+}$).
The continual Scheme 1 also minimizes the current loss to convergence, i.e., until it is perfectly solved (in the realizable case). In fact, Evron et al. (2022) identified the following reduction.
@@
In the realizable case (Assumption 1) under any ordering $\tau$, continual linear regression learned to convergence333 The continual learner minimizes $\mathcal{L}_{\tau(t)}$ with (S)GD to convergence; they do not explicitly compute pseudo-inverses. is equivalent to the block Kaczmarz method. That is, the iterates $\mathbf{w}_{0},\dots,\mathbf{w}_{k}$ of Schemes 1 and 2 coincide.

### v2

@@ Abstract
We theoretically study the common continual learning setup where an overparameterized model is sequentially fitted to a set of jointly realizable tasks. We analyze the forgetting—loss on previously seen tasks—after $k$ iterations. For continual linear models, we prove that fitting a task is equivalent to a single stochastic gradient descent (SGD) step on a modified objective. We develop novel last-iterate SGD upper bounds in the realizable least squares setup, which we then leverage to derive new results for continual learning. Focusing on random orderings over $T$ tasks, we establish universal forgetting rates, whereas existing rates depend on the problem dimensionality or complexity. Specifically, in continual regression with replacement, we improve the best existing rate from $\mathcal{O}((d-\bar{r})/k)$ to $\mathcal{O}(\min(1/\sqrt[4]{k},\sqrt{d-\bar{r}}/k,\sqrt{T\bar{r}}/k))$, where $d$ is the dimensionality and $\bar{r}$ the average task rank. Furthermore, we establish the first rate for random task orderings without replacement. The obtained rate of $\mathcal{O}(\min(1/\sqrt[4]{T},\,(d-r)/T))$ proves for the first time that randomization alone—with no task repetition—can prevent catastrophic forgetting in sufficiently long task sequences. Finally, we prove a matching $\mathcal{O}(1/\sqrt[4]{k})$ forgetting rate for continual linear classification on separable data. Our universal rates apply for broader projection methods, such as block Kaczmarz and POCS, illuminating their loss convergence under i.i.d. and one-pass orderings.
@@ 1 Introduction
In continual learning (CL), tasks are presented sequentially, one at a time. The goal is for the learner to adapt to the current task—e.g., by fine-tuning using gradient-based algorithms—while retaining knowledge from previous tasks. A central challenge in this setting is termed catastrophic forgetting, where expertise from earlier tasks is lost when adapting to newer ones. Forgetting is influenced by factors such as task similarity and overparameterization [20], and is also related to trade-offs like the plasticity-stability dilemma [46]. CL is becoming increasingly important with the rise of foundation models, where retraining is prohibitively expensive and data from prior tasks is often unavailable, e.g., due to privacy or data retention constraints.
@@
Previous work has shown, both analytically [15, 16, 34, 31, 9] and empirically [39, 26], that forgetting diminishes over time when task ordering is cyclic or random. Different orderings can be explored from multiple perspectives: as a strategy to mitigate forgetting (e.g., by actively ordering an agent’s learning environments); as a naturally occurring phenomenon, such as periodic trends in e-commerce; or as a means to model and analyze popular CL benchmarks, such as randomly split datasets.
@@
Our work focuses on a widely studied analytical setting—realizable continual linear regression, [^1] Despite its simplicity, the continual linear regression setting is insightful enough to investigate different factors and aspects in continual settings, e.g., task similarity [15, 42, 27], task recurrence [15, 34], overparameterization [20], and algorithms [13, 55]. We follow prior theoretical work assuming tasks are jointly realizable [e.g., 15, 34, 20]. Other notable work alternatively allow label noise, at the cost of assuming either i.i.d. features [19, 42, 2] or commutative covariance matrices [40, 41, 72]—while our analysis facilitates any data matrices. where $T$ tasks are learned sequentially over $k$ iterations in a uniform random ordering. Evron et al. [15] established that the worst-case expected forgetting lies between $\Omega\left({{1}/{k}}\right)$ and $\mathcal{O}\left({{(d-\bar{r})}/{k}}\right)$, where $d$ is the problem dimensionality, and $\bar{r}$ the average rank of individual data matrices. This raises a fundamental question, critical in highly overparameterized regimes: Does worst-case forgetting necessarily scale with dimensionality, and if so, is the dependence indeed linear?
@@
To this end, we bridge continual learning and the literature on last-iterate stochastic gradient descent (SGD) analysis. We revisit an established connection between continual linear regression and the Kaczmarz method for solving systems of linear equations [32, 15]. Given rank-$1$ tasks, this method is known to perform a normalized stochastic gradient step on the least squares objective, fully minimizing the current task’s loss and implying a “stepwise-optimal” step size. Deepening this connection, we prove that even for general data ranks, learning a task in continual linear regression and performing an update in the Kaczmarz method, are both equivalent to a single SGD step on a modified objective with a constant, stepwise-optimal step size.
@@
Motivated by this, we prove convergence rates for the last iterate of fixed-step size SGD that, crucially, hold for a broad range of step sizes not covered by prior work [e.g., 62, 18, 4, 73, 67, 69]. Curiously, prior results either hold only for the average iterate [e.g., 1], or for small step sizes, bounded away from the stepwise-optimal step size crucial for the continual learning setup [e.g., 67]. We overcome this challenge through a careful combination of analysis techniques for SGD [65, 62], further tightening the analysis to accommodate a wider range of step sizes, including the stepwise-optimal one.
@@
Applying our last-iterate analysis to continual regression, we tighten the existing forgetting rate and establish the first dimension-independent rate (see Table 1). Furthermore, we provide the first rate for random task orderings without replacement, proving that task repetition is not obligatory to guarantee convergence when ${k=T\to\infty}$, thus highlighting the effect of randomization as compared to repetition. Our results also yield novel rates for the closely related Kaczmarz and NLMS methods.
@@
Finally, we prove a matching rate for the squared loss of the broader Projection Onto Convex Sets framework [22]. This extends our results to continual linear classification on separable data—previously linked to projection algorithms [16]—and provides this setting’s first universal rate, independent of the problem’s “complexity”.
@@
To summarize, our main contributions in this paper are:
  • We establish a new reduction from continual linear regression to SGD with a particular choice of a “stepwise-optimal” step size, generalizing ideas from prior work that only applied to rank-$1$ tasks, to tasks of arbitrary rank. This facilitates last-iterate analysis for studying forgetting.
@@
  • Our main contribution, building on these techniques, is a set of improved rates of forgetting in continual linear regression, including the first universal rates, independent of the problem dimensionality or complexity, as well as the first rates for without-replacement orderings, indicating task repetition is not mandatory to diminish forgetting. See Table 1 for a summary.
@@ 2 Main Setting: Continual Linear Regression
We mainly investigate the fundamental continual linear regression setting, as studied in many theoretical papers. This setting is easy-to-analyze, yet often sheds light on important CL phenomena.1
@@
Formally, the learner is given a collection of $T$ linear regression tasks, $(\mathbf{X}_{1},\mathbf{y}_{1}),\dots,(\mathbf{X}_{T},\mathbf{y}_{T})$, where $\mathbf{X}_{m}\!\in\!\mathbb{R}^{n_{m}\!\times d},\mathbf{y}_{m}\!\in\!\mathbb{R}^{n_{m}}$. For $k$ iterations, tasks are learned according to a task ordering $\tau:\left[{k}\right]\!\to\!\left[{T}\right]$. We analyze random orderings, previously studied in continual linear models [e.g., 15, 16, 31].
Definition 2.1 (Random Task Ordering).
@@
We are now ready to define the learning scheme we study, which, at each iteration, naively minimizes the sum of squared errors for the current regression task.222 This objective is natural for regression; our analysis also extends to the mean squared error (refining our $R$).
@@
   For each iteration $t=1,\dots,k$:
       $\mathbf{w}_{t}\leftarrow$ Start from $\mathbf{w}_{t-1}$ and minimize the current task’s loss $\mathcal{L}_{\tau(t)}(\mathbf{w})\triangleq{\frac{1}{2}}\left\|{\mathbf{X}_{\tau({t})}\mathbf{w}-\mathbf{y}_{\tau({t})}}\right\|^{2}$           with (S)GD to convergence [^3] Learning to convergence facilitates the analysis, but other analytical choices exist [see 31].
@@
This scheme was previously linked to the Kaczmarz method and, in a special case, to normalized SGD [15]. In Section 3, we explain and develop these connections to enable novel analysis.
Our main assumption is the existence of offline solutions that perfectly solve all $T$ tasks jointly, as assumed in much of the theoretical CL literature [e.g., 15, 16, 20, 31]. This assumption simplifies the analysis1 and rules out cases where forgetting previous tasks is beneficial, as new tasks may directly contradict them. Finally, this assumption is reasonable in highly overparameterized models and is thus linked to the linear dynamics of deep networks in the neural tangent kernel (NTK) regime [29].
Assumption 2.2 (Joint Linear Realizability).
@@
Definition 2.3 (Minimum-Norm Offline Solution).
We denote, $\mathbf{w}_{\star}\triangleq{\operatorname*{argmin}}_{\mathbf{w}\in\mathcal{W}_{\star}}\left\|{\mathbf{w}}\right\|$.
Commonly in continual learning setups, the model performance on past tasks degrades, sometimes significantly, even in linear models [15]. Our goal is to bound this degradation, i.e., “forgetting". Following common definitions [e.g., 13, 16], we define forgetting as the average increase in the loss of the last iterate on previous tasks.
Definition 2.4 (Forgetting).
@@
Under arbitrary orderings, Evron et al. [15] showed forgetting can be “catastrophic” in the sense that $\lim_{k\to\infty}\mathbb{E}\left[F_{\tau}(k)\right]>0$. However, as we show, this cannot happen under the random ordering.
Remark 2.5 (Forgetting vs. Regret).
While regret and forgetting are related, they can differ significantly [15]. Regret is a key quantity in online learning, defined as $\frac{1}{2k}{\sum}_{t=1}^{k}\|{\mathbf{X}_{\tau(t)}\mathbf{w}_{t-1}-\mathbf{y}_{\tau(t)}}\|^{2}$ in our setting. That is, it measures the suboptimality of each iterate on the consecutive task. In contrast, forgetting evaluates an iterate’s performance across earlier tasks.
@@
Definition 2.6 (Training Loss).
@@
We bound both the forgetting and the loss, leveraging a key property—expected (in-sample) forgetting can be upper bounded using expected training loss across all tasks. Specifically, Lemma B.1 (in App. B) states that $\mathbb{E}_{\tau}[F_{\tau}(k)]\leq 2\mathbb{E}_{\tau}\left[\mathcal{L}\left(\mathbf{w}_{k-1}\right)\right]+\frac{\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}}{k}$ in orderings with replacement, where $R\triangleq\max_{m\in\left[{T}\right]}\left\|{\mathbf{X}_{m}}\right\|$ is the data “radius” and the dependence of $\mathbf{w}_{k-1}$ on $\tau_{1},\dots,\tau_{k-1}$ is implicit. Without-replacement orderings yield a related but more refined bound. The additive $\frac{\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}}{k}$ term is negligible compared to other terms in our bounds.
@@ 3 Reductions: From Continual Linear Regression to Kaczmarz to SGD
Prior work has drawn connections between continual linear regression and the Kaczmarz method [15], which we revisit pedagogically to keep the paper self-contained. Importantly, this leads to novel links between CL, the Kaczmarz method, and SGD on special functions (Schemes 1,2,3), allowing us to improve the rates for continual and Kaczmarz methods by analyzing the last iterate of SGD instead.
@@
  Input: ${f_{m}(\mathbf{w})\!=\!\frac{1}{2}\!\left\|{\mathbf{X}_{m}^{+}\left({\mathbf{X}_{m}\mathbf{w}\!-\!\mathbf{y}_{m}}\right)}\right\|^{2}\!,\forall m\!\in\!\left[{T}\right]}$
@@ 3.1 Revisit: Continual Linear Regression and the Kaczmarz Method
The (block) Kaczmarz method in Scheme 2 [32, 14] is a classical iterative method for solving a linear system $\mathbf{X}\mathbf{w}=\mathbf{y}$, easily mapped to our learning problem by stacking tasks in blocks, i.e.,
[…]
In each iteration, the Kaczmarz method (Scheme 2) perfectly solves the current block, i.e., ${\mathbf{X}_{\tau(t)}\mathbf{w}_{t}=\mathbf{y}_{\tau(t)}}$ (to see that, recall that $\mathbf{X}_{\tau(t)}^{+}$ denotes the Moore-Penrose pseudo-inverse of $\mathbf{X}_{\tau(t)}^{+}$).
The continual Scheme 1 also minimizes the current loss to convergence, i.e., until it is perfectly solved (in the realizable case). In fact, Evron et al. [15] identified the following reduction.
@@
In the realizable case (Assumption 2.2) under any ordering $\tau$, continual linear regression learned to convergence444 The learner minimizes $\mathcal{L}_{\tau(t)}$ with (S)GD to convergence; the pseudo-inverse is not computed explicitly. is equivalent to the block Kaczmarz method. That is, the iterates $\mathbf{w}_{0},\dots,\mathbf{w}_{k}$ of Schemes 1 and 2 coincide.